Exercise 5.60 Revie A sample of gas has a mass of 0.555 8. Its volume is 115 mL at a temperature of 87 °C and a pressure of 7 circle answers
Answer #1

Given:

P = 756.0 torr

= (756.0/760) atm

= 0.9947 atm

V = 115.0 mL

= (115.0/1000) L

= 0.115 L

T = 87.0 oC

= (87.0+273) K

= 360 K

find number of moles using:

P * V = n*R*T

0.9947 atm * 0.115 L = n * 0.08206 atm.L/mol.K * 360 K

n = 3.872*10^-3 mol

mass(solute)= 0.555 g

use:

number of mol = mass / molar mass

3.872*10^-3 mol = (0.555 g)/molar mass

molar mass = 1.433*10^2 g/mol

Answer: 143 g/mol